Two dice are rolled 20 times. Compute Pe(13), the empirical probability of obtaining a score of 13. Explain the result very briefly.
Answer by venugopalramana(3286)

Two dice are rolled 20 times. Compute Pe(13), the empirical probability of obtaining a score of 13. Explain the result very briefly.
TOTAL NO.OF POSSIBILITIES =6*6=36
13 CAN NOT OCCURR IN 2 DICE THROW AS MAXIMUM SCORE IS 6+6=12
SO P=0
PE=20*0=0